# Levels of AGI: Operationalizing Progress on the Path to AGI

### v2

@@ 4 Levels of AGI
wide range of non-physical tasks, including metacognitive abilities like learning new skills
Level 0: No AI
Narrow Non-AI
@@
ChatGPT (OpenAI, 2023), Bard (Anil et al., 2023), Llama 2 (Touvron et al., 2023), Gemini (Pichai and Hassabis, 2023)
@@
Table 1: A leveled, matrixed approach toward classifying systems on the path to AGI based on depth (performance) and breadth (generality) of capabilities. Example systems in each cell are approximations based on current descriptions in the literature or experiences interacting with deployed systems. Unambiguous classification of AI systems will require a standardized benchmark of tasks, as we discuss in the Testing for AGI section. Note that general systems that broadly perform at a level N may be able to perform a narrow subset of tasks at higher levels. The "Competent AGI" level, which has not been achieved by any public systems at the time of writing, best corresponds to many prior conceptions of AGI, and may precipitate rapid social change once achieved.
In accordance with Principle 2 ("Focus on Generality and Performance") and Principle 6 ("Focus on the Path to AGI, not a Single Endpoint"), in Table 1 we introduce a matrixed leveling system that focuses on performance and generality as the two dimensions that are core to AGI:
@@
While this taxonomy rates systems according to their performance, systems that are capable of achieving a certain level of performance (e.g., against a given benchmark) may not match this level in practice when deployed. For instance, user interface limitations may reduce deployed performance. Consider the example of DALLE-2 (Ramesh et al., 2022), which we estimate as a Level 3 Narrow AI (“Expert Narrow AI”) in our taxonomy. We estimate the “Expert” level of performance since DALLE-2 produces images of higher quality than most people are able to draw; however, the system has failure modes (e.g., drawing hands with incorrect numbers of digits, rendering nonsensical or illegible text) that prevent it from achieving a “Virtuoso” performance designation. While theoretically an “Expert” level system, in practice the system may only be “Competent,” because prompting interfaces are too complex for most end-users to elicit optimal performance (as evidenced by user studies (Zamfirescu-Pereira et al., 2023) and by the existence of marketplaces (e.g., PromptBase ) in which skilled prompt engineers sell prompts). This observation emphasizes the importance of designing ecologically valid benchmarks (that would measure deployed rather than idealized performance) as well as the importance of considering how human-AI interaction paradigms interact with the notion of AGI (a topic we return to in the “Capabilities vs. Autonomy” Section).
The highest level in our matrix in terms of combined performance and generality is ASI (Artificial Superintelligence). We define "Superhuman" performance as outperforming 100% of humans. For instance, we posit that AlphaFold (Jumper et al., 2021; Varadi et al., 2021) is a Level 5 Narrow AI ("Superhuman Narrow AI") since it performs a single task (predicting a protein’s 3D structure from an amino acid sequence) above the level of the world’s top scientists. This definition means that Level 5 General AI ("ASI") systems will be able to do a wide range of tasks at a level that no human can match. Additionally, this framing also implies that Superhuman systems may be able to perform an even broader generality of tasks than lower levels of AGI, since the ability to execute tasks that qualitatively differ from existing human skills would by definition outperform all humans (who fundamentally cannot do such tasks). For example, non-human skills that an ASI might have could include capabilities such as neural interfaces (perhaps through mechanisms such as analyzing brain signals to decode thoughts (Tang et al., 2023; Bellier et al., 2023)), oracular abilities (perhaps through mechanisms such as analyzing large volumes of data to make high-quality predictions (Schoenegger and Park, 2023)), or the ability to communicate with animals (perhaps by mechanisms such as analyzing patterns in their vocalizations, brain waves, or body language (Goldwasser et al., 2023; Andreas et al., 2022)).
@@ 5 Testing for AGI
Our intent is that an AGI benchmark would include a broad suite of cognitive and metacognitive tasks (per Principle 3), measuring diverse properties including (but not limited to) linguistic intelligence, mathematical and logical reasoning (Webb et al., 2023), spatial reasoning, interpersonal and intra-personal social intelligences, the ability to learn new skills (Chollet, 2019), and creativity. A benchmark might include tests covering psychometric categories proposed by theories of intelligence from psychology, neuroscience, cognitive science, and education; however, such “traditional” tests must first be evaluated for suitability for benchmarking computing systems, since many may lack ecological and construct validity in this context (Serapio-García et al., 2023).
One open question for benchmarking performance is whether to allow the use of tools, including potentially AI-powered tools, as an aid to human performance. This choice may ultimately be task dependent and should account for ecological validity in benchmark choice (per Principle 5). For example, in determining whether a self-driving car is sufficiently safe, benchmarking against a person driving without the benefit of any modern AI-assisted safety tools would not be the most informative comparison; since the relevant counterfactual involves some driver-assistance technology, we may prefer a comparison to that baseline.
@@
Determining that something is not an AGI at a given level simply requires identifying several [^4] We hesitate to specify the precise number or percentage of tasks that a system must pass at a given level of performance in order to be declared a General AI at that Level (e.g., a rule such as ”a system must pass at least 90% of an AGI benchmark at a given performance level to get that rating”). While we think this will be a very high percentage, it will probably not be 100%, since it seems clear that broad but imperfect generality is impactful (individual humans also lack consistent performance across all possible tasks, but remain generally intelligent). Determining what portion of benchmarking tasks at a given level demonstrate generality remains an open research question. tasks that people can typically do but the system cannot adequately perform. Systems that pass the majority of the envisioned AGI benchmark at a particular performance level ("Emerging," "Competent," etc.), including new tasks added by the testers, can be assumed to have the associated level of generality for practical purposes (i.e., though in theory there could still be a test the AGI would fail, at some point unprobed failures are so specialized or atypical as to be practically irrelevant).
Developing an AGI benchmark will be a challenging and iterative process. It is nonetheless a valuable north-star goal for the AI research community. Measurement of complex concepts may be imperfect, but the act of measurement helps us crisply define our goals and provides an indicator of progress.
6 Risk in Context: Autonomy and Human-AI Interaction
Discussions of AGI often include discussion of risk, including "x-risk" – existential (for AI Safety, 2023) or other very extreme risks (Shevlane et al., 2023). A leveled approach to defining AGI enables a more nuanced discussion of how different combinations of performance and generality relate to different types of AI risk. While there is value in considering extreme risk scenarios, understanding AGI via our proposed ontology rather than as a single endpoint (per Principle 6) can help ensure that policymakers also identify and prioritize risks in the near-term and on the path to AGI.
@@ 6.1 Levels of AGI as a Framework for Risk Assessment
As we advance along our capability levels toward ASI, new risks are introduced, including misuse risks, alignment risks, and structural risks (Zwetsloot and Dafoe, 2019). For example, the “Expert AGI” level is likely to involve structural risks related to economic disruption and job displacement, as more and more industries reach the substitution threshold for machine intelligence in lieu of human labor. On the other hand, reaching “Expert AGI” likely alleviates some risks introduced by “Emerging AGI” and “Competent AGI,” such as the risk of incorrect task execution. The “Virtuoso AGI” and “ASI” levels are where many concerns relating to x-risk are most likely to emerge (e.g., an AI that can outperform its human operators on a broad range of tasks might deceive them to achieve a mis-specified goal, as in misalignment thought experiments (Christian, 2020)).
@@
We acknowledge that whether an AGI benchmark should include tests for potentially dangerous capabilities (e.g., the ability to deceive, to persuade (Veerabadran et al., 2023), or to perform advanced biochemistry (Morris, 2023)) is controversial. We lean on the side of including such capabilities in benchmarking, since most such skills tend to be dual use (having valid applications to socially positive scenarios as well as nefarious ones). Dangerous capability benchmarking can be de-risked via Principle 4 (Potential, not Deployment) by ensuring benchmarks for any dangerous or dual-use tasks are appropriately sandboxed and not defined in terms of deployment. However, including such tests in a public benchmark may allow malicious actors to optimize for these abilities; understanding how to mitigate risks associated with benchmarking dual-use abilities remains an important area for research by AI safety, AI ethics, and AI governance experts.
Concurrent with this work, Anthropic released Version 1.0 of its Responsible Scaling Policy (RSP) (Anthropic, 2023b). This policy uses a levels-based approach (inspired by biosafety level standards) to define the level of risk associated with an AI system, identifying what dangerous capabilities may be associated with each AI Safety Level (ASL), and what containment or deployment measures should be taken at each level. Current SOTA generative AIs are classified as an ASL-2 risk. Including items matched to ASL capabilities in any AGI benchmark would connect points in our AGI taxonomy to specific risks and mitigations.
@@
While capabilities provide prerequisites for AI risks, AI systems (including AGI systems) do not and will not operate in a vacuum. Rather, AI systems are deployed with particular interfaces and used to achieve particular tasks in specific scenarios. These contextual attributes (interface, task, scenario, end-user) have substantial bearing on risk profiles. AGI capabilities alone do not determine destiny with regards to risk, but must be considered in combination with contextual details.
@@
Using an AI system to advance scientific discovery (e.g., protein-folding)
Possible:
@@
Table 2: More capable AI systems unlock new human-AI interaction paradigms (including fully autonomous AI). The choice of appropriate autonomy level need not be the maximum achievable given the capabilities of the underlying model. One consideration in the choice of autonomy level are resulting risks. This table’s examples illustrate the importance of carefully considering human-AI interaction design decisions.
@@
These Levels of Autonomy are correlated with the Levels of AGI. Higher levels of autonomy are “unlocked” by AGI capability progression, though lower levels of autonomy may be desirable for particular tasks and contexts (including for safety reasons) even as we reach higher levels of AGI. Carefully considered choices around human-AI interaction are vital to safe and responsible deployment of frontier AI models.
@@
Unlike prior taxonomies of computer automation proposed by Sheridan and colleagues (Sheridan et al., 1978; Sheridan and Parasuraman, 2005; Parasuraman et al., 2000) that take a computer-centric perspective (framing automation in terms of how much control the designer relinquishes to computers), we characterize the concept of autonomy through the lens of the nature of human-AI interaction style; further, our ontology considers how AI capabilities may enable particular interaction paradigms and how the combination of level of autonomy and level of AGI may impact risk. Shneiderman (Shneiderman, 2020) observes that automation is not a zero-sum game, and that high levels of automation can co-exist with high levels of human control; this view is compatible with our perspective of considering automation through the perspective of varying styles of human-AI partnerships.
@@
We emphasize the importance of the “No AI” paradigm. There may be many situations where this is desirable, including for education, enjoyment, assessment, or safety reasons. For example, in the domain of self-driving vehicles, when Level 5 Self-Driving technology is widely available, there may be reasons for using a Level 0 (No Automation) vehicle. These include for instructing a new driver (education), for pleasure by driving enthusiasts (enjoyment), for driver’s licensing exams (assessment), or in conditions where sensors cannot be relied upon such as technology failures or extreme weather events (safety). While Level 5 Self-Driving (SAE International, 2021) vehicles would likely be a Level 5 Narrow AI (“Superhuman Narrow AI”) under our taxonomy [^5] Fully autonomous vehicles might arguably be classified as Level 4 Narrow AI (”Virtuoso Narrow AI”) per our taxonomy; however, we suspect that in practice autonomous vehicles may need to reach the Superhuman performance standard to achieve widespread social acceptance regarding perceptions of safety, illustrating the importance of contextual considerations., the same considerations regarding human vs. computer autonomy apply to AGIs. We may develop an AGI, but choose not to deploy it autonomously (or choose to deploy it with differentiated autonomy levels in distinct circumstances as dictated by contextual considerations).
@@
Certain aspects of generality may be required to make particular interaction paradigms desirable. For example, the Autonomy Levels 3, 4, and 5 ("Collaborator," "Expert," and "Agent") may only work well if an AI system also demonstrates strong performance on certain metacognitive abilities (learning when to ask a human for help, theory of mind modeling, social-emotional skills). Implicit in our definition of Autonomy Level 5 ("AI as an Agent") is that such a fully autonomous AI can act in an aligned fashion without continuous human oversight, but knows when to consult humans (Shah et al., 2021). Interfaces that support human-AI alignment through better task specification, the bridging of process gulfs, and evaluation of outputs (Terry et al., 2023) are a vital area of research for ensuring that the field of human-computer interaction keeps pace with the challenges and opportunities of interacting with AGI systems.
6.2.1 Human-AI Interaction Paradigm as a Framework for Risk Assessment

### v3

@@ 4 Levels of AGI
Table 1: A leveled, matrixed approach toward classifying systems on the path to AGI based on depth (performance) and breadth (generality) of capabilities. The assignment of example systems to cells is approximate. Unambiguous classification of AI systems will require a standardized benchmark of tasks, as we discuss in Section 5. Note that general systems that broadly perform at a level N may be able to perform a narrow subset of tasks at higher levels. The “Competent AGI” level, which has not been achieved by any public systems at the time of writing, best corresponds to many prior conceptions of AGI, and may precipitate rapid societal change once achieved.
@@
wide range of non-physical tasks, including metacognitive tasks like learning new skills
Level 0: No AI Narrow Non-AI
@@
ChatGPT (OpenAI, 2023), Bard (Anil et al., 2023), Llama 2 (Touvron et al., 2023), Gemini (Pichai & Hassabis, 2023)
@@
In accordance with Principle 2 (“Focus on Generality and Performance”) and Principle 6 (“Focus on the Path to AGI, not a Single Endpoint”), in Table 1 we introduce a matrixed leveling system that focuses on performance and generality as the two dimensions that are core to AGI:
@@
While this taxonomy rates systems according to their performance, systems that are capable of achieving a certain level of performance (e.g., against a given benchmark) may not match this level in practice when deployed. For instance, user interface limitations may reduce deployed performance. Consider DALLE-2 (Ramesh et al., 2022), which we estimate as a Level 3 Narrow AI (“Expert Narrow AI”) in our taxonomy. We estimate the “Expert” level of performance since DALLE-2 produces images of higher quality than most people are able to draw; however, the system has failure modes (e.g., drawing hands with incorrect numbers of digits, rendering nonsensical or illegible text) that prevent it from achieving a “Virtuoso” performance designation. While theoretically an “Expert” level system, in practice the system may only be “Competent,” because prompting interfaces are too complex for most end-users to elicit optimal performance (as evidenced by user studies (Zamfirescu-Pereira et al., 2023) and the existence of marketplaces (e.g., PromptBase ) in which skilled prompt engineers sell prompts). This observation emphasizes the importance of designing ecologically valid benchmarks (that approximate deployed rather than idealized performance), as well as the importance of considering the human-AI interaction paradigms.
The highest level in our matrix in terms of combined performance and generality is ASI (Artificial Superintelligence). We define “Superhuman” performance as outperforming 100% of humans. For instance, we posit that AlphaFold (Jumper et al., 2021; Varadi et al., 2021) is a Level 5 Narrow AI (“Superhuman Narrow AI”) since it performs a single task (predicting a protein’s 3D structure from an amino acid sequence) above the level of the world’s top scientists. This definition means that Level 5 General AI (“ASI”) systems will be able to do a wide range of tasks at a level that no human can match. Additionally, this framing also implies that Superhuman systems may be able to perform an even broader generality of tasks than lower levels of AGI, since the ability to execute tasks that qualitatively differ from existing human skills would by definition outperform all humans (who fundamentally cannot do such tasks). For example, non-human skills that an ASI might have could include capabilities such as neural interfaces (perhaps through mechanisms such as analyzing brain signals to decode thoughts (Tang et al., 2023; Bellier et al., 2023)), oracular abilities (perhaps through mechanisms such as analyzing large volumes of data to make high-quality predictions (Schoenegger & Park, 2023)), or the ability to communicate with animals (perhaps by mechanisms such as analyzing patterns in their vocalizations, brain waves, or body language (Goldwasser et al., 2023; Andreas et al., 2022)).
@@ 5 Testing for AGI
Our intent is that an AGI benchmark would include a broad suite of cognitive and metacognitive tasks (per Principle 3), measuring diverse properties including (but not limited to) linguistic intelligence, mathematical and logical reasoning (Webb et al., 2023), spatial reasoning, interpersonal and intra-personal social intelligences, the ability to learn new skills (Chollet, 2019), and creativity. A benchmark might include tests covering psychometric categories proposed by theories of intelligence from psychology, neuroscience, cognitive science, and education; however, such tests must first be evaluated for suitability for benchmarking computing systems, since many may lack ecological and construct validity in this context (Serapio-García et al., 2023).
We emphasize the importance of metacognition, and suggest that an AGI benchmark should include metacognitive tasks such as (1) the ability to learn new skills, (2) the ability to know when to ask for help, and (3) social metacognitive abilities such as those relating to theory of mind. The ability to learn new skills (Chollet, 2019) is essential to generality, since it is infeasible for a system to be optimized for all possible use cases a priori; this necessitates related sub-skills such as the ability to select appropriate strategies for learning (Pressley et al., 1987). Knowing when to ask for help is necessary to support alignment and appropriate human-AI interaction (Terry et al., 2023), and would include an awareness of the limits of the model’s own abilities (Demetriou & Kazi, 2006), which relates to the sub-skill of model calibration (Liang et al., 2023), i.e., the model’s ability to proactively anticipate and retroactively evaluate how well it would do/did on certain tasks. Additionally, theory of mind tasks are sometimes considered metacognitive (Tullis & Fraundorf, 2017), though are sometimes classified separately as social cognition (Gardner, 2011); the ability of systems to accurately model end-users is a necessary component of alignment for AGI systems.
One open question for benchmark design is whether to allow the use of tools, including potentially AI-powered tools, as an aid to human performance. This choice may ultimately be task dependent and should account for ecological validity in benchmark choice (per Principle 5). For example, in determining whether a self-driving car is sufficiently safe, benchmarking against a person driving without the benefit of any modern AI-assisted safety tools would not be the most informative comparison; since the relevant counterfactual involves some driver-assistance technology, we may prefer a comparison to that baseline.
@@
Determining that something is not an AGI at a given level simply requires identifying tasks that people can typically do but the system cannot adequately perform. Systems that pass the majority of the envisioned AGI benchmark at a particular performance level (“Emerging,” “Competent,” etc.), including new tasks added by the testers, can be assumed to have the associated level of generality for practical purposes (i.e., though in theory there could still be a test the AGI would fail, at some point unprobed failures are so specialized or atypical as to be practically irrelevant). We hesitate to specify the number or percentage of tasks that a system must pass at a given level of performance in order to be declared a General AI at that Level (e.g., a rule such as “a system must pass at least 90% of an AGI benchmark at a given performance level to get that rating”). While we think this will be a very high percentage, it will probably not be 100%, since it seems clear that broad but imperfect generality is impactful (individual humans also lack consistent performance across all possible tasks, but are generally intelligent). Determining what portion of benchmarking tasks at a given level demonstrate generality remains an open research question.
6 Risk, Autonomy, and Interaction
Discussions of AGI often include discussion of risk, including “x-risk” – existential (for AI Safety, 2023) or other very extreme risks (Shevlane et al., 2023). A leveled approach to defining AGI enables a more nuanced discussion of how different combinations of performance and generality relate to different types of AI risk. While there is value in considering extreme risk scenarios, understanding AGI via our proposed ontology rather than as a single endpoint (per Principle 6) can help ensure that policymakers also identify and prioritize risks in the near-term and on the path to AGI.
@@ 6.1 Levels of AGI as a Framework for Risk Assessment
As we advance along our capability levels toward ASI, new risks are introduced, including misuse risks, alignment risks, and structural risks (Zwetsloot & Dafoe, 2019). For example, the “Expert AGI” level is likely to involve structural risks related to economic disruption and job displacement, as more and more industries reach the substitution threshold for machine intelligence in lieu of human labor. On the other hand, reaching “Expert AGI” likely alleviates some risks introduced by “Emerging AGI” and “Competent AGI,” such as the risk of incorrect task execution. The “Virtuoso AGI” and “ASI” levels are where many concerns relating to x-risk are most likely to emerge (e.g., an AI that can outperform its human operators on a broad range of tasks might deceive them to achieve a mis-specified goal, as in misalignment thought experiments (Christian, 2020)).
@@
Whether an AGI benchmark should include tests for potentially dangerous capabilities (e.g., the ability to deceive, to persuade (Veerabadran et al., 2023), or to perform advanced biochemistry (Morris, 2023)) is controversial. We lean on the side of including such capabilities in benchmarking, since most such skills tend to be dual use (having valid applications to socially positive scenarios as well as nefarious ones). Dangerous capability benchmarking can be de-risked via Principle 4 (Potential, not Deployment) by ensuring benchmarks for any dangerous or dual-use tasks are appropriately sandboxed and not defined in terms of deployment. However, including such tests in a public benchmark may allow malicious actors to optimize for these abilities; understanding how to mitigate risks associated with benchmarking dual-use abilities remains an important area for research by AI safety, AI ethics, and AI governance experts.
Concurrent with this work, Anthropic released Version 1.0 of its Responsible Scaling Policy (RSP) (Anthropic, 2023b). This policy uses a levels-based approach (inspired by biosafety levels (Richmond & McKinney, 2009)) to define the level of risk associated with an AI system, identifying what dangerous capabilities may be associated with each AI Safety Level (ASL), and what containment or deployment measures should be taken at each level. Current SOTA generative AIs are classified as an ASL-2 risk. Including items matched to ASL capabilities in any AGI benchmark would connect points in our AGI taxonomy to specific risks and mitigations.
@@
Table 2: More capable AI systems unlock new human-AI interaction paradigms. The choice of appropriate autonomy level need not be the maximum achievable given the capabilities of the underlying model. One consideration in the choice of autonomy level are resulting risks. This table’s examples illustrate the importance of carefully considering human-AI interaction design decisions.
@@
Using an AI system to advance scientific discovery (e.g., protein-folding) Possible:
@@
While capabilities provide prerequisites for AI risks, AI systems (including AGI systems) do not and will not operate in a vacuum. Rather, AI systems are deployed with particular interfaces and used to achieve particular tasks in specific scenarios. These contextual attributes (interface, task, scenario, end-user) have substantial bearing on risk.
@@
These Levels of Autonomy are correlated with the Levels of AGI. Higher levels of autonomy are “unlocked” by AGI capability progression, though lower levels of autonomy may be desirable for particular tasks and contexts even as we reach higher levels of AGI. Carefully considered choices around human-AI interaction are vital to safe and responsible deployment of frontier AI models.
@@
Unlike prior taxonomies of computer automation (Sheridan et al., 1978; Sheridan & Parasuraman, 2005; Parasuraman et al., 2000) that take a computer-centric perspective (framing automation in terms of how much control the designer relinquishes to computers), we characterize the concept of autonomy through the lens of the nature of human-AI interaction style; further, our ontology considers how AI capabilities may enable particular interaction paradigms and how the combination of level of autonomy and level of AGI may impact risk. Shneiderman (Shneiderman, 2020) observes that automation is not a zero-sum game, and that high levels of automation can co-exist with high levels of human control; this view is compatible with our perspective of considering automation through the perspective of varying styles of human-AI partnerships.
@@
We emphasize the importance of the “No AI” paradigm for many contexts, including for education, enjoyment, assessment, or safety reasons. For example, in the domain of self-driving vehicles, when Level 5 Self-Driving technology is widely available, there may be reasons for using a Level 0 (No Automation) vehicle. These include for instructing a new driver (education), for pleasure by driving enthusiasts (enjoyment), for driver’s licensing exams (assessment), or in conditions where sensors cannot be relied upon such as technology failures or extreme weather events (safety). While Level 5 Self-Driving (SAE International, 2021) vehicles would likely be a Level 4 or 5 Narrow AI under our taxonomy, the same considerations regarding human vs. computer autonomy apply to AGIs. We may develop an AGI, but choose not to deploy it autonomously, or choose to deploy it with differentiated autonomy levels in distinct circumstances as dictated by contextual considerations.
@@
Certain aspects of generality may be required to make particular interaction paradigms desirable. For example, the Autonomy Levels 3, 4, and 5 (“Collaborator,” “Expert,” and “Agent”) may only work well if an AI system also demonstrates strong performance on certain metacognitive abilities (learning when to ask a human for help, theory of mind modeling, social-emotional skills). Implicit in our definition of Autonomy Level 5 (“AI as an Agent”) is that such a fully autonomous AI can act in an aligned fashion without continuous human oversight, but knows when to consult humans (Shah et al., 2021). Interfaces that support human-AI alignment through better task specification, the bridging of process gulfs, and evaluation of outputs (Terry et al., 2023) are a vital area of research.
6.3 Human-AI Interaction and Risk Assessment